OpenStudy (anonymous):

How can i find the margin of sampling error? P=16%, n=1000

OpenStudy (anonymous):

I dont understand how im able to find it ive tried \[\sqrt{p(p-1)/n}\] but i never get any of the choices

OpenStudy (kropot72):

You need to be given the critical value of z for a specific confidence level. Is a value of confidence percentage stated in the question. Perhaps you should post the complete question

OpenStudy (anonymous):

The whole question is "Find the margin of sampling error to the nearest percent." P=16%, n=1000

OpenStudy (kropot72):

The formula that you have used is for the standard error, and is not for the margin of error.

OpenStudy (mathmale):

Did you attempt to evaluate that expression on a calculator? If so, try putting down the expression onto paper and doing the work on paper as a check. I would write \[\sqrt{p(p-1)/n}~as~\sqrt{\frac{ p(1-p) }{ n }}\]

OpenStudy (mathmale):

Note how you have p-1 and I have 1-p? That alone may explain why none of your answers have been correct.

OpenStudy (kropot72):

The margin of error is the product of the critical z-value and the standard error.\[\large Margin\ of\ Error=z _{c}\sqrt{\frac{p(1-p)}{n}}\] where zc is the critical value of z.

OpenStudy (anonymous):

Ok so i need to find the z value before i can find the margin of error?

OpenStudy (kropot72):

You can find the critical value of z when you know the required confidence level. This must be known. Perhaps your question is part of a larger question?

OpenStudy (anonymous):

I wrote everything it gave me I'm not really sure.

OpenStudy (kropot72):

Well it is not possible to answer this question unless the required confidence level is known.

OpenStudy (anonymous):

oh ok ill talk to my teacher about it tomorrow. Thanks for the help anyway though guys.

OpenStudy (kropot72):

For a 95% confidence level, the answer would be 0.0227 (or 2.27%)

OpenStudy (kropot72):

What are the answer choices?

OpenStudy (anonymous):

A. 2% B. 4% C. .016% D. .04%

OpenStudy (kropot72):

If the confidence level was 90%, the margin of error is 1.9%.

OpenStudy (kropot72):

Choices C and D are unrealistic. The most likely choice is A.

OpenStudy (mathmale):

Good idea to look up "z critical value" and learn what it represents and how to find its specific value for a given level of confidence. \[z _{c}=1.96 \]

OpenStudy (mathmale):

applies for the 95% level of confidence, 1.645 for the 90% level, and so on. These are often shown in tables of z-scores, and can be found easily on a TI-83 or -84 calculator, once you've done this a few times.